OpenStudy (anonymous):

Evaluate the definite integral. Use a graphing utility to verify your result.

OpenStudy (anonymous):

\[\int\limits_{-1}^{0} (t^\frac{ 1 }{ 3 }-t^\frac{ 2 }{ 3 }) dt\]

OpenStudy (anonymous):

\[\frac{3}{40} \left(9+i \sqrt{3}\right) \]